please show me how to work out the problem 144=-12(x+5)
Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

144=-12(x+5)
Divide thru by 12
12 = -(x+5)
Divide by -1
-12 = x+5
Subtract 5
x = -17

Answer by Theo(13342)

144 = -12 * (x + 5)

Remove the parentheses by multiplying each term within the parentheses by -12 to get:

144 = (-12 * x) + (-12 * 5)

Simplify this by multiplying out the terms and combining like terms to get:

144 = -12*x - 60

add 60 to both sides of this equation to get:

144 + 60 = -12*x

combine like terms to get:

204 = -12*x

divide both sides of this equation by (-12) to get:

204 / (-12) = (-12) * x / (-12)

Simplify and combine like terms to get:

-17 = x

You have x = -17 as your answer.

replace this value for x in the original equation as shown below:

original equation is 144 = -12 * (x + 5)

replace x with -17 to get:

144 = -12 * (-17 + 5)

combine like terms within parentheses to get:

144 = -12 * (-12)

simplify and combine like terms to get:

144 = 144